# On the complexity of estimating ground state entanglement and free energy

### v1

@@ 1 Introduction
Given is a $k$-local Hamiltonian $H_{AB}=\sum_{j=1}^{m}H_{j}$ on registers $A$ and $B$ of $n_{A}+n_{B}=n$ qubits, energy threshold parameters $\alpha,\beta$ such that $\beta-\alpha=1/\operatorname{poly}(n)$ and entropy threshold parameters $s,t$ such that $s-t=1/\operatorname{poly}(n)$. Decide:
@@
  • (NO case) If $\forall\;\left|\psi\right\rangle_{AB}$ either $\left\langle\psi\right|H\left|\psi\right\rangle\geq\beta$ or $S(\psi_{A})\leq t$.
@@
$\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-complete for $k\geq 5$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
@@
We tie the complexity of our next two problems to $\mathsf{QMA}(2)$ [KMY03], a class notorious for resisting anything but non-trivial complexity bounds $\mathsf{QMA}\subseteq\mathsf{QMA}(2)\subseteq\mathsf{NEXP}$.
@@
  • (NO case) If $\forall\;\left|\psi\right\rangle_{AB}$ either $\left\langle\psi\right|H\left|\psi\right\rangle\geq\beta$ or $S(\rho_{A})\geq s$.
@@
For $k\geq 5$, $\mathsf{LELES}$ is $\mathsf{QMA}(2)$-hard, even when $s-t=\Omega(n)$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
Thus, upper bounding the complexity of LELES is at least as difficult as bounding the power of $\mathsf{QMA}(2)$, a longstanding open problem. Whether this implies finding low entanglement ground states is easier (LELES) or harder than finding high entanglement ground states (HELES) depends on what one believes regarding the power of $\mathsf{QMA}$ versus $\mathsf{QMA}(2)$. It is entirely plausible that $\mathsf{QMA}(2)=\mathsf{QMA}\subsetneq\mathsf{qq\mhyphen QAM}$ (i.e. LELES is easier than HELES), or that $\mathsf{QMA}(2)=\mathsf{NEXP}\supsetneq\mathsf{qq\mhyphen QAM}$ (i.e. HELES is much harder than LELES).
@@
(Low Energy Approximate Product State ($\mathsf{LEAPS}$)) Given is a $k$-local Hamiltonian $H$ on $n$ qubits, energy thresholds $\alpha,\beta$, and distance thresholds $a,b<1$. Decide:
@@
$\mathsf{LEAPS}$ is $\mathsf{QMA}$-complete if $\beta-\alpha-a\|H\|_{\infty}\geq\frac{1}{\operatorname{poly}(n)}$ and $b>a$.
@@
We make use of two main technical tools, which we encapsulate into two standalone lemmas (Lemma 3.1 and Lemma 3.2, respectively) for ease of further adoption. The first is a $\mathsf{qq\mhyphen QAM}$ protocol for entropy verification, which we use to prove that $\mathsf{qq\mhyphen QAM}$ contains $\mathsf{HELES}$ and $\mathsf{FEA}$. This protocol is not new; it appears implicitly in [KGN19, proof of Lemma 33] and is based on ideas from [BST08, Section 5]. The second tool is a channel-to-Hamiltonian construction, which we use to prove that $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard and to prove that $\mathsf{LELES}$ and $\mathsf{LEAPS}$ are $\mathsf{QMA}(2)$-hard. We now sketch each of these tools.
@@ 2.1 Complexity classes and problems
The input to a $\mathsf{CIMM}$ instance consists of a quantum channel $\Phi$ and constants $a,b\in(0,1)$ such that $(1-a)^{2}>1-b^{2}$. The task is to distinguish the following two cases:
@@
The input to a $\mathsf{MaxOutQEA}$ instance consists of a quantum channel $\Phi$ and a parameter $t\in\mathbb{R}$. The task is to distinguish the following two cases:
@@
The following $\mathsf{QMA}(2)$-complete problem will be very useful for our purposes.
@@ 2.2 Useful definitions and lemmas
For all $n,k,\varepsilon$, there exist explicit $(k,d,\varepsilon)$-quantum extractor $T\colon L((\mathbb{C}^{2})^{\otimes n})\to L((\mathbb{C}^{2})^{\otimes n})$ with $d=2(n-k+2\log\frac{1}{\varepsilon})+O(1)$, or $d=n-k+2\log\frac{1}{\varepsilon}+2\log n+O(1)$, depending on the construction used.
@@
Let $\rho_{A},\sigma_{A}$ be quantum states such that $\lVert\rho_{A}-\sigma_{A}\rVert_{\mathrm{tr}}\leq\varepsilon$ and let $\left|\phi_{\rho}\right\rangle_{AB},\left|\phi_{\sigma}\right\rangle_{AB}$ be purifications of $\rho,\sigma$. Then, there exists a unitary $U_{B}$ acting only on the purifying register $B$ such that $F\left(\lvert\phi_{\sigma}\rangle\langle\phi_{\sigma}\rvert,(I_{A}\otimes U_{B})\lvert\phi_{\rho}\rangle\langle\phi_{\rho}\rvert(I_{A}\otimes U_{B})\right)\geq(1-2\varepsilon)^{2}$.
@@ 3.1 Entropy verification in $\mathsf{qq\mhyphen QAM}$
    Output the resulting state $\sigma_{\mathcal{A}\mathcal{B}}$.
@@ 3.2 Channel-to-Hamiltonian construction
We will use Kitaev’s circuit-to-Hamiltonian construction applied to the unitary $U_{\Phi}$ with two modification. We use no output penalty and post idle for $L$ time steps. The Hamiltonian $H_{\Phi}$ will be given by
@@ 4 The High Entropy Ground State problem
We will show the slightly stronger statement that $\mathsf{HELES}\in\mathsf{qq\mhyphen QAM}$, even if we replace the pure states in the definition of $\mathsf{HELES}$ with mixed states. [^12] Note that the YES and NO cases of the pure state variant are subsets of the YES and NO cases of the mixed state variant. Hence the mixed state variant is at least as hard as the pure state variant. In fact, our results show that they are equally hard.
@@
For $k\geq 5$, $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard.
@@
We show a poly-time many-one reduction from $\mathsf{MaxOutQEA}$. Let $(\Phi,\tau)$ be a $\mathsf{MaxOutQEA}$ instance. Without loss of generality we can assume that for all $\sigma$, there is a pure state $\left|\psi\right\rangle$ with $\Phi(\sigma)=\Phi(\psi)$. [^13] To see this, modify $\Phi\colon L(\mathcal{H}_{A})\to L(\mathcal{H}_{B})$ by adding another copy of the input register that is immediately traced out. Taking $\left|\psi\right\rangle$ to be some purification of $\sigma$ then suffices. Let $H_{\Phi}$ be the Hamiltonian acting on $A,B,C,E$ resulting from Lemma 3.2, with $L$ to be determined later. The output of the reduction will be $(H_{\Phi},\alpha,\beta,s,t)$, where $\alpha,\beta,s$ and $t$ will also be determined later.
@@
For soundness, suppose that for all $\sigma_{A}$, $S\left(\Phi(\sigma)\right)\leq\tau$ and let $\left|\psi\right\rangle_{ABCE}$ be such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\beta$ ($\beta$ will be determined later). By property 2 of Lemma 3.2, there is a pure state $\left|\phi\right\rangle_{A}$ such that
$T\coloneq\|\psi_{B}-\Phi(\phi_{A})\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}.$ (70)
@@
we see that $S(\psi_{B})\leq t$ as desired. Note that Eqs. 69 and 73 can be satisfied by an $L$ that is polynomially large in $n=n_{A}+n_{B}$. ∎
@@ 5 Free energy and Gibbs states
    The verifier runs the entropy verification protocol with $\delta=\frac{b-a}{4\|H\|_{\infty}},\delta^{\prime}<(1+\frac{1}{\beta})^{-1}\frac{b-a}{4}$ and $\tau=\tilde{S}$. They reject if the protocol rejects.
@@ 6 Low Entropy Ground State Problem
We will reduce from the $\mathsf{QMA}(2)$-complete problem $\mathsf{PPIO}$ (Definition 2.7). Let $(U,a,b)$ be a $\mathsf{PPIO}$ instance and assume that $a=2^{-\Omega(n)}$ and $b=2-2^{-\Omega(n)}$. We map this to a $\mathsf{LELES}$ instance $(H_{U},\alpha,\beta,s,t)$ where $H_{U}$ is the Hamiltonian acting on registers $A,B,C$ resulting from the circuit-to-Hamiltonian construction (Lemma 3.2) applied to the channel $\Phi(\rho)\coloneq U(\rho_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert)U^{\dagger}$. Note that $H_{\Phi}$ is indeed $5$-local. $s$ and $t$ will be $\Omega(n)$ and $O(1)$ respectively and the parameters $\alpha,\beta$ will be fixed at a later point in the proof.
@@
Suppose now that for all $\left|\psi\right\rangle$, $\Phi(\psi)$ is $b$ far from any product state. Let $\left|\psi\right\rangle_{ABC}$ be such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\beta$. By the second property in Lemma 3.2 there exists a pure state $\left|\chi\right\rangle_{A}$ such that
$T\coloneq\|\psi_{AB}-\Phi(\chi)\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}.$ (89)
@@
$\geq\Omega(n)-\left(2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\left(n-\log\left(2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}\right)\right).$ (91)
Setting $\beta=\Delta/L\geq\Omega(L^{-4})$, and choosing $L$ to be a sufficiently large polynomial in $n$ we get $S(\psi_{B})\geq\Omega(n)$ as desired, completing the proof.
@@ 7 A $\mathsf{QMA}(2)$-complete local Hamiltonian problem
$|\operatorname{Tr}(H\rho)-\operatorname{Tr}(H\sigma)|\leq\|H(\rho-\sigma)\|_{1}\leq\|H\|_{\infty}\|\rho-\sigma\|_{1}.$ (92)
Hence, if $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$ and $\|\psi-\phi_{L}\otimes\phi_{R}\|_{1}\leq a$, then $\left\langle\phi_{R}\right|\left\langle\phi_{L}\right|H\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle\leq\alpha+a\|H\|_{\infty}$. To solve the $\mathsf{LEAPS}$ instance $(H,\alpha,\beta,a,b)$ we can now solve the $\mathsf{SeparableLocalHamiltonian}$ instance $(H,\alpha^{\prime},\beta)$ where $\alpha^{\prime}=\alpha+a\|H\|_{\infty}$. Note that $\beta-\alpha^{\prime}\geq\frac{1}{\operatorname{poly}(n)}$ by assumption. Furthermore, note that if $(H,\alpha,\beta,a,b)$ is a YES instance of $\mathsf{LEAPS}$, then $(H,\alpha^{\prime},\beta)$ is a YES instance of $\mathsf{SeparableLocalHamiltonian}$. If $(H,\alpha,\beta,a,b)$ is a NO instance, then certainly all product states have energy $\geq\beta$, hence $(H,\alpha^{\prime},\beta)$ is a NO instance as well. Note that we do not need a gap between $a$ and $b$ as this subsumed by the energy gap.
For hardness we reduce from the Separable Local Hamiltonian problem. Map an instance $(H,\alpha,\beta)$ to the $\mathsf{LEAPS}$ instance $(H,\alpha,\beta^{\prime},0,b)$, where $b\leq\frac{\beta-\alpha}{2\|H\|_{\infty}}$ and $\beta^{\prime}=\beta-\|H\|_{\infty}b$. Note that $\beta^{\prime}-\alpha\geq\frac{\beta-\alpha}{2}\geq\frac{1}{\operatorname{poly}(n)}$. Completeness follows trivially. For soundness, let $\left|\psi\right\rangle$ be a state that is $b$ close to a product state $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$. By the Hölder inequality, $\left\langle\psi\right|H\left|\psi\right\rangle\geq\left\langle\phi_{R}\right|\left\langle\phi_{L}\right|H\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle-\|H\|_{\infty}b\geq\beta-\|H\|_{\infty}b=\beta^{\prime}$, completing the proof. ∎
@@
We begin with the characterization $\mathsf{QMA}(k)=\mathsf{QMA}(2)$ from [HM13]. The proofs will be $\left|\psi_{LR}\right\rangle,\left|\phi_{L}\right\rangle$ and $\left|\phi_{R}\right\rangle$. With probability $1/2$ each, the verifier performs one of the following checks:

### v2

@@ 1 Introduction
Given is a $k$-local Hamiltonian $H_{AB}=\sum_{j=1}^{m}H_{j}$ on registers $A$ and $B$ of $n_{A}+n_{B}=n$ qubits with norm $\|H\|_{\infty}\leq\operatorname{poly}(n)$, energy threshold parameters $\alpha,\beta$ such that $\beta-\alpha=1/\operatorname{poly}(n)$ and entropy threshold parameters $s,t$ such that $s-t=1/\operatorname{poly}(n)$. Decide:
@@
  • (NO case) If $\forall\;\rho_{AB}$ either $\operatorname{Tr}(H\rho_{AB})\geq\beta$ or $S(\rho_{A})\leq t$.
@@
$\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-complete for $k\geq 5$. Containment holds for all $n_{A},n_{B}$ and hardness requires $n_{B}=\operatorname{poly}(n_{A})$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
@@
We tie the complexity of our next two problems to $\mathsf{QMA}(2)$ [KMY03], a class notorious for resisting anything but the trivial complexity bounds $\mathsf{QMA}\subseteq\mathsf{QMA}(2)\subseteq\mathsf{NEXP}$.
@@
  • (NO case) If $\forall\;\rho_{AB}$ either $\operatorname{Tr}(H\rho)\geq\beta$ or $S(\rho_{A})\geq s$.
@@
For $k\geq 5$, $\mathsf{LELES}$ is $\mathsf{QMA}(2)$-hard, even when $s-t=\Omega(n)$. We require $n_{B}=\operatorname{poly}(n_{A})$. Hardness remains even when restricting to physically motivated Hamiltonians such as the 2D Heisenberg model (Corollary 8.6).
Thus, upper bounding the complexity of LELES is at least as difficult as bounding the power of $\mathsf{QMA}(2)$, a longstanding open problem. Whether this implies finding low entanglement ground states is easier (LELES) or harder than finding high entanglement ground states (HELES) depends on what one believes regarding the power of $\mathsf{QMA}$ versus $\mathsf{QMA}(2)$. It is entirely plausible that $\mathsf{QMA}(2)=\mathsf{QMA}\subsetneq\mathsf{qq\mhyphen QAM}$ (i.e. LELES is easier than HELES), or that $\mathsf{QMA}(2)=\mathsf{NEXP}\supsetneq\mathsf{qq\mhyphen QAM}$ (i.e. LELES is much harder than HELES).
@@
(Low Energy Approximate Product State ($\mathsf{LEAPS}$)) Given is an $n$ qubit $k$-local Hamiltonian $H=\sum_{i=1}^{m}H_{i}$ with $\|H_{i}\|_{\infty}\leq 1$ for all $i$, energy thresholds $\alpha,\beta$, and distance thresholds $a,b<1$. Decide:
@@
Let $d$ be an efficiently computable upper bound on $\|H\|_{\infty}$ such as the number of local terms $m$. Then $\mathsf{LEAPS}$ is $\mathsf{QMA}$-complete if $\beta-\alpha-ad\geq\frac{1}{\operatorname{poly}(n)}$ and $b>a$.
@@
We make use of two main technical tools, which we encapsulate into two standalone lemmas (Lemma 3.1 and Lemma 3.2, respectively) for ease of further adoption. The first is a $\mathsf{qq\mhyphen QAM}$ protocol for entropy verification, which we use to prove that $\mathsf{qq\mhyphen QAM}$ contains $\mathsf{HELES}$ and $\mathsf{FEA}$. This protocol is not new; it appears implicitly in [KGN19, proof of Lemma 33] and is based on ideas from [BST08, Section 5]. The second tool is a “channel-to-Hamiltonian” construction: Kitaev’s circuit-to-Hamiltonian construction applied to the Stinespring representation of a channel with minor modifications. We use this to prove that $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard and to prove that $\mathsf{LELES}$ and $\mathsf{LEAPS}$ are $\mathsf{QMA}(2)$-hard. We now sketch each of these tools.
@@ 2.1 Complexity classes and problems
The input to a $\mathsf{CIMM}$ instance consists of a quantum channel $\Phi$ (given as a circuit implementing its Stinespring representation) and constants $a,b\in(0,1)$ such that $(1-a)^{2}>1-b^{2}$. The task is to distinguish the following two cases:
@@
The input to a $\mathsf{MaxOutQEA}$ instance consists of (the Stinespring representation of) a quantum channel $\Phi$ and a parameter $t\in\mathbb{R}$. The task is to distinguish the following two cases:
@@
The following $\mathsf{QMA}(2)$-complete problem due to [GHMW13] will be very useful for our purposes.
@@ 2.2 Useful definitions and lemmas
For all $n,k,\varepsilon$, there exist constructions of explicit $(k,d,\varepsilon)$-quantum extractors $T\colon L((\mathbb{C}^{2})^{\otimes n})\to L((\mathbb{C}^{2})^{\otimes n})$ with $d=2(n-k+2\log\frac{1}{\varepsilon})+O(1)$, or $d=n-k+2\log\frac{1}{\varepsilon}+2\log n+O(1)$, depending on the construction used.
@@
Let $\rho_{A},\sigma_{A}$ be quantum states such that $\|\rho_{A}-\sigma_{A}\|_{1}\leq\varepsilon$ and let $\left|\phi_{\rho}\right\rangle_{AB},\left|\phi_{\sigma}\right\rangle_{AB}$ be purifications of $\rho,\sigma$. Then, there exists a unitary $U_{B}$ acting only on the purifying register $B$ such that $F\left(\lvert\phi_{\sigma}\rangle\langle\phi_{\sigma}\rvert,(I_{A}\otimes U_{B})\lvert\phi_{\rho}\rangle\langle\phi_{\rho}\rvert(I_{A}\otimes U_{B})\right)\geq(1-2\varepsilon)^{2}$.
@@ 3.1 Entropy verification in $\mathsf{qq\mhyphen QAM}$
    Accept and output the resulting state $\sigma_{\mathcal{A}\mathcal{B}}$.
@@ 3.2 Channel-to-Hamiltonian construction
We will use Kitaev’s circuit-to-Hamiltonian construction applied to the unitary $U_{\Phi}$ with two modifications. We use no output penalty and post idle for $L$ time steps. The Hamiltonian $H_{\Phi}$ will be given by
@@ 4 The High Entropy Ground State problem
We will show the slightly stronger statement that $\mathsf{HELES}\in\mathsf{qq\mhyphen QAM}$, even if we replace the pure state in the YES case of the definition of $\mathsf{HELES}$ with a mixed state. [^12] Note that the YES case of the pure state variant is a subset of the YES case of the mixed state variant. Hence the mixed state variant is at least as hard as the pure state variant. In fact, our results show that they are equally hard.
@@
For $k\geq 5$ and $n_{B}=\operatorname{poly}(n_{A})$, $\mathsf{HELES}$ is $\mathsf{qq\mhyphen QAM}$-hard.
@@
We show a poly-time many-one reduction from $\mathsf{MaxOutQEA}$. Let $(\Phi,\tau)$ be a $\mathsf{MaxOutQEA}$ instance. Without loss of generality we can assume that for all $\sigma$, there is a pure state $\left|\psi\right\rangle$ with $\Phi(\sigma)=\Phi(\psi)$. [^13] To see this, modify $\Phi\colon L(\mathcal{H}_{A})\to L(\mathcal{H}_{B})$ by adding another copy of the input register that is immediately traced out. Taking $\left|\psi\right\rangle$ to be some purification of $\sigma$ then suffices. Let $H_{\Phi}$ be the Hamiltonian acting on $A,B,C,E$ resulting from Lemma 3.2, with $L$ to be determined later. The output of the reduction will be $(H_{\Phi},\alpha,\beta,s,t)$, where the cut is between the $B$ and $ACE$ registers. The parameters $\alpha,\beta,s$ and $t$ will be determined later.
@@
For soundness, suppose that for all $\sigma_{A}$, $S\left(\Phi(\sigma)\right)\leq\tau$ and let $\rho_{ABCE}$ be such that $\operatorname{Tr}(H\rho_{ABCE})\leq\beta$ ($\beta$ will be determined later). By property 2 of Lemma 3.2, there is a state $\sigma_{A}$ such that
$T\coloneq\|\rho_{B}-\Phi(\sigma_{A})\|_{1}\leq 2\sqrt{\frac{\beta}{\Delta}}+\frac{2T}{T+L+1}.$ (70)
@@
we see that $S(\rho_{B})\leq t$ as desired. Note that Eqs. 69 and 73 can be satisfied by an $L$ that is polynomially large in $n=n_{A}+n_{B}$.
Close inspection of [KGN19] shows that $\mathsf{qq\mhyphen QAM}$-hard $\mathsf{MaxOutQEA}$ instances can be assumed to map $n$ qubits to $n$ qubits. It follows that the size of the $ACE$ register is polynomially larger than that of the $B$ register. ∎
@@ 5 Free energy and Gibbs states
    The verifier runs the entropy verification protocol with [^14] While the spectral norm of $H$ might be unknown to the verifier, they can trivially compute an upper bound. $\delta\leq\frac{b-a}{4\|H\|_{\infty}},\delta^{\prime}<(1+\frac{1}{\beta})^{-1}\frac{b-a}{4}$ and $\tau=\tilde{S}$. They reject if the protocol rejects.
@@ 6 Low Entropy Ground State Problem
We will reduce from the $\mathsf{QMA}(2)$-complete problem $\mathsf{PPIO}$ (Definition 2.7). Let $(U,a,b)$ be a $\mathsf{PPIO}$ instance and assume that $a=2^{-\Omega(n)}$ and $b=2-2^{-\Omega(n)}$. We map this to a $\mathsf{LELES}$ instance $(H_{U},\alpha,\beta,s,t)$ where $H_{U}$ is the Hamiltonian acting on registers $A,B,C$ resulting from the circuit-to-Hamiltonian construction (Lemma 3.2) applied to the channel $\Phi(\rho)\coloneq U(\rho_{A}\otimes\lvert 0^{n_{B}}\rangle\langle 0^{n_{B}}\rvert)U^{\dagger}$. The cut will be between the $B$ and $AC$ registers. Note that $H_{\Phi}$ is indeed $5$-local. The parameters $s$ and $t$ will be $\Omega(n)$ and $O(1)$ respectively and $\alpha,\beta$ will be fixed at a later point in the proof.
@@
Suppose now that for all $\left|\psi\right\rangle$, $\Phi(\psi)$ is $b$ far from any product state. We will first consider pure states: let $\left|\psi\right\rangle_{ABC}$ be such that $\left\langle\psi\right|H\left|\psi\right\rangle\leq\beta^{\prime}$. By the second property in Lemma 3.2 there exists a pure state $\left|\chi\right\rangle_{A}$ such that
$T\coloneq\|\psi_{AB}-\Phi(\chi)\|_{1}\leq 2\sqrt{\frac{\beta^{\prime}}{\Delta}}+\frac{2T}{T+L+1}.$ (89)
@@
$\geq\Omega(n)-\left(2\sqrt{\frac{\beta^{\prime}}{\Delta}}+\frac{2T}{T+L+1}\right)\left(n_{B}-\log\left(2\sqrt{\frac{\beta^{\prime}}{\Delta}}+\frac{2T}{T+L+1}\right)\right).$ (91)
Setting $\beta^{\prime}=\Delta/L\geq\Omega(L^{-4})$, and choosing $L$ to be a sufficiently large polynomial in $n$ we get $S(\psi_{B})\geq\Omega(n)$ as desired.
We now turn to mixed states. We have already shown that all low energy pure states have high entropy. Let $\rho_{ABC}$ be such that $\operatorname{Tr}(H\rho_{ABC})\leq\beta=\frac{\beta^{\prime}}{n}$. We can write $\rho_{ABC}$ as a convex combination of pure states. All pure states with energy $\leq\beta^{\prime}$ will have entropy $\Omega(n)$, and the weight on pure states with energy $\beta^{\prime}$ cannot be more than $\frac{1}{n}$ without breaking the energy constraint. By concavity of the entropy, we then have $S(\rho_{B})\geq\frac{n-1}{n}\Omega(n)=\Omega(n)$, completing the proof.
Close inspection of [GHMW13] shows that a $\mathsf{QMA}(2)$-hard $\mathsf{PPIO}$ instance can be assumed to map $n$ qubits to $\Theta(n)$ qubits. It follows that the size of $AC$ is polynomial in the size of $B$.
@@ 7 A $\mathsf{QMA}(2)$-complete local Hamiltonian problem
$|\operatorname{Tr}(H\rho)-\operatorname{Tr}(H\sigma)|\leq\|H(\rho-\sigma)\|_{1}\leq\|H\|_{\infty}\|\rho-\sigma\|_{1}\leq d\|\rho-\sigma\|_{1}.$ (92)
Hence, if $\left\langle\psi\right|H\left|\psi\right\rangle\leq\alpha$ and $\|\psi-\phi_{L}\otimes\phi_{R}\|_{1}\leq a$, then $\left\langle\phi_{R}\right|\left\langle\phi_{L}\right|H\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle\leq\alpha+ad$. To solve the $\mathsf{LEAPS}$ instance $(H,\alpha,\beta,a,b)$ we can now solve the $\mathsf{SeparableLocalHamiltonian}$ instance $(H,\alpha^{\prime},\beta)$ where $\alpha^{\prime}=\alpha+ad\geq\alpha+a\|H\|_{\infty}$. Note that $\beta-\alpha^{\prime}\geq\frac{1}{\operatorname{poly}(n)}$ by assumption. Furthermore, note that if $(H,\alpha,\beta,a,b)$ is a YES instance of $\mathsf{LEAPS}$, then $(H,\alpha^{\prime},\beta)$ is a YES instance of $\mathsf{SeparableLocalHamiltonian}$. If $(H,\alpha,\beta,a,b)$ is a NO instance, then certainly all product states have energy $\geq\beta$, hence $(H,\alpha^{\prime},\beta)$ is a NO instance as well. Note that we do not need a gap between $a$ and $b$ as this subsumed by the energy gap.
For hardness we reduce from the Separable Local Hamiltonian problem. Map an instance $(H,\alpha,\beta)$ to the $\mathsf{LEAPS}$ instance $(H,\alpha,\beta^{\prime},0,b)$, where $b=\frac{\beta-\alpha}{2d}\leq\frac{\beta-\alpha}{2\|H\|_{\infty}}$ and $\beta^{\prime}=\beta-bd\leq\beta-\|H\|_{\infty}b$. Note that $\beta^{\prime}-\alpha\geq\frac{\beta-\alpha}{2}\geq\frac{1}{\operatorname{poly}(n)}$. Completeness follows trivially. For soundness, let $\left|\psi\right\rangle$ be a state that is $b$ close to a product state $\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle$. By the Hölder inequality, $\left\langle\psi\right|H\left|\psi\right\rangle\geq\left\langle\phi_{R}\right|\left\langle\phi_{L}\right|H\left|\phi_{L}\right\rangle\left|\phi_{R}\right\rangle-\|H\|_{\infty}b\geq\beta-db=\beta^{\prime}$, completing the proof. ∎
@@
We begin with the characterization $\mathsf{QMA}(k)=\mathsf{QMA}(2)$ from [HM13]. The proofs will be $\left|\psi_{LR}\right\rangle,\left|\phi_{L}\right\rangle$ and $\left|\phi_{R}\right\rangle$. With probability $1/2$ each, the verifier performs one of the following checks. They accept iff the performed check accepts.